OpenStudy (anonymous):

5. Use the functions to answer the question. f(x) = 4x – 1; g(x) = 5x^2 Which is equal to (f circle g) (x)? (Points : 4) 80x^2 – 40x + 5 20x^3 – 5x^2 5x^2 + 4x – 1 20x^2 – 1

OpenStudy (anonymous):

\[f\circ g(x)\] means \[f(g(x))\]

OpenStudy (anonymous):

step one is to write \[f\circ g(x)=f(g(x))=f(5x^2)\]

OpenStudy (anonymous):

then since \[f(\diamondsuit)=4\diamondsuit -1\] you have \[f(5x^2)=4\times 5x^2-1\]